OpenStudy (anonymous):

Which numbers are solutions of the equation? |5 – 3x| – 3 = 2

OpenStudy (uri):

|5 – 3x| – 3 = 2 5-3x-3=2 2-2=3x 0=3x |5 – 3x| – 3 = 2 5-3x-3=-2 2-2=3x 0=3x No solution! :p

OpenStudy (anonymous):

fax you ya smart cookie

OpenStudy (anonymous):

but that's not the correct choice my friend ;P

OpenStudy (uri):

Yeah hold on lol

OpenStudy (uri):

|5 – 3x| – 3 = 2 5-3x-3=-2 2+2=3x 4=3x x=4/3

OpenStudy (anonymous):

WEEEEEEEEEEE

OpenStudy (anonymous):

\(|5-3x| - 3 = 2\) Case 1: \(x\le \frac{5}{3}\) \(\Rightarrow |5-3x|\) = \(5 - 3x\) \(|5-3x| - 3 = 2\) \(\Leftrightarrow 5 - 3x -3=2\) \(\Leftrightarrow 3x=0\) ---> no solution Case 2: \(x>\frac{5}{3}\) \(\Rightarrow |5-3x|\) = \(-5 + 3x\) \(|5-3x| - 3 = 2\) \(\Leftrightarrow -5 + 3x -3=2\) \(\Leftrightarrow 3x = 10\) \(\Leftrightarrow x= \frac{10}{3}\) (satisfy)